OpenStudy (anonymous):

What is 6 times 4/40

OpenStudy (triciaal):



OpenStudy (triciaal):

then simplify by canceling common factors from top and bottom

jhonyy9 (jhonyy9):

4/40 = 1/10 6*1/10 = ? hope this will help you